# Think-to-Talk or Talk-to-Think? When LLMs Come Up with an Answer in Multi-Hop Arithmetic Reasoning

### v3

@@ Abstract
This study investigates the incremental, internal problem-solving process of language models (LMs) with arithmetic multi-hop reasoning as a case study. We specifically investigate when LMs internally resolve sub/whole problems through first reading the problem statements, generating reasoning chains, and achieving the final answer to mechanistically interpret LMs’ multi-hop problem-solving process. Our experiments reveal a systematic incremental reasoning strategy underlying LMs. They have not derived an answer at the moment they first read the problem; instead, they obtain (sub)answers while generating the reasoning chain. Therefore, the generated reasoning chains can be regarded as faithful reflections of the model’s internal computation.
[…]
100
Think-to-Talk or Talk-to-Think?
When LLMs Come Up with an Answer in Multi-Hop Arithmetic Reasoning
Keito Kudo1,2, Yoichi Aoki1,2, Tatsuki Kuribayashi3, Shusaku Sone1 Masaya Taniguchi2, Ana Brassard2,1, Keisuke Sakaguchi1,2, Kentaro Inui3,1,2 1Tohoku University, 2RIKEN, 3MBZUAI {keito.kudo.q4, youichi.aoki.p2, sone.shusaku.r8}@dc.tohoku.ac.jp, {tatsuki.kuribayashi, kentaro.inui}@mbzuai.ac.ae, keisuke.sakaguchi@tohoku.ac.jp, {masaya.taniguchi, ana.brassard} @riken.jp
@@ 1 Introduction
An explanation may be produced in two modes: as a post hoc explanation to a predetermined conclusion (Think-to-Talk) or by the process of reaching a conclusion while explaining (Talk-to-Think). An analogy applies to large language models (LLMs) using chain-of-thought (CoT; Wu et al., 2023a) reasoning: is generated CoT reasoning chain a post-hoc explanation, or does it reflect real-time step-by-step solving? This question is particularly relevant to the (mechanistic) interpretablity of LLMs — how models incrementally resolve the multi-hop problem, in other words, how information internally flows to reach the final decision.
In this study, we identify such internal reasoning patterns of LLMs. There may be multiple possible internal strategies for LLMs under the CoT scenario, i.e., first feed the problem statements and then have models generate full reasoning chains. One presumably hard strategy for LLMs would be to reach the final answer even during the first pass of problem statements before CoT generation, and then while generating CoT reasoning chains, the model simply refers to their predetermined answers as a post-hoc explanation (think-to-talk mode). The opposite will be that models do nothing during the first pass of the problem statements and start solving the problem after CoT generation begins (talk-to-think mode). Such different internal mechanisms can not be distinguished by just observing the model outputs; rather, one has to interpret and intervene in their internal representations during their multi-hop CoT-style reasoning.
@@
Figure 1: Using linear probes, we investigated at which time during the LLM’s problem-solving process it is possible to determine the values of each variable, illustrating the model’s problem-solving process. Our analysis indicates that LMs come up with (sub)answers during CoT (second pass). This conclusion is also consistent with the findings from the causal experiments in §˜5.
@@
In our first experiments, we apply linear probes to model internals at each layer at each timestep to determine when answers are reached internally. We prepared controlled testbeds of symbolic arithmetic reasoning tasks and observed whether trained probes could accurately predict and control the values of specific variables (Figure 1). By comparing accuracies across each timestep, one can observe at which point models internally start being informative to the probes, illustrating the model’s internal reasoning flow.
The results reveal common patterns across models. They have not derived an answer at the moment they first read the problem (first-pass); instead, they obtain (sub)answers while generating the reasoning chain. These tendencies are consistent and systematic across our different testbeds.
Based on the above finding, we further conducted causal intervention analyses to clarify the causal relationship between the model’s internal representations and the final answer (§˜5). We found that, when generating the reasoning chain, the model’s internal problem-solving process exhibits a recency bias, relying heavily on the information in the immediately preceding portion of the chain. In other words, the generated reasoning chains can be regarded as faithful reflections of the model’s internal computation.  [^1] Code and data will be made available upon acceptance.
@@
Table 1: Examples of arithmetic reasoning tasks used in our experiments at each complexity level. #Step indicates the number of required operations to reach the final answer. #Stack indicates how many variables’ values are not immediately determined in their first appearing equation. #Dist. is the number of unnecessary distractor equations. The number (e.g., $\mathbf{-3}$) indicated in the lower right corner of each equation represents the equation’s position. This position is used as a reference point for calculating $t^{*}_{\mathrm{eq}}$ in §˜4.2.
2 Related Work
Multi-stage human incremental language comprehension.
Humans sometimes make several attempts with different strategies or mindsets, particularly when resolving a complex task. A common view may be, for example, that humans first adopt a shallow, fast solution, and once it fails, switch to a more expensive, presumably accurate one. Such a multi-stage processing is even related to the recent debate on the cognitive plausibility of LLMs in computational psycholinguistics Oh and Schuler (2023); Shain et al. (2024); Kuribayashi et al. (2024); Gruteke Klein et al. (2024). LLMs have now been criticized because they can not estimate human cognitive costs incurred, possibly by the multi-stage nature of human sentence processing (i.e., reanalysis) van Schijndel and Linzen (2021); Huang et al. (2024); humans re-read the sentence again from earlier points as an additional trial when facing difficulties in comprehension, but LLMs do not explicitly have such a multi-stage mechanism. Although this study is not focused on sentence processing or LMs’ humanlikeness, one critical question in this line is whether and how LLMs switch their reasoning strategy for the same problem through multiple trials, and how to track their dynamic internal states. In our case, we simplify the setting to forced-decode the common chain-of-thought style format as two-time trials of the problem, where an LM first reads the problem statement and then (re-)analyzes the problem while generating the reasoning chain as well as copying the problem statement again. Here, we analyze what kind of process LLMs perform, particularly in the first pass of the problem statement, and how their initial processing is related to their later generation of reasoning chain and answer. Our results suggest some surprising behaviors; models systematically come up with answers to simple subproblems in the first pass, but these computations are not reused in the second pass of CoT-style reasoning, suggesting some redundancy in their internal reasoning.
Interpreting multi-hop reasoning in language models.
Interpreting internal mechanisms of LMs has been actively investigated Conneau et al. (2018); Tenney et al. (2019); Niven and Kao (2019); nostalgebraist (2020); Geva et al. (2023); Lieberum et al. (2024); Ghandeharioun et al. (2024); Ferrando and Voita (2024). They revealed, for example, specialized attention heads responsible for specific operations (Cabannes et al., 2024) or decision-making, copying, and induction (Dutta et al., 2024). With a more concrete example, Yang et al. (2024c) showed that, even during the first pass of the roblem statements such as The mother of the singer of Thriller is ___, language models first resolve a bridge entity, Stevie Wonder in this case, then identify the final answer. This study is more focused on the difference between the first pass of the problem statements (before CoT generation) and their second pass involving explicit problem solving (while CoT generation).
Arithmetic representations in LLMs.
How models handle numerical information has also been closely studied. For instance, Heinzerling and Inui (2024) used partial least squares regression Wold et al. (2001) to demonstrate that numeric attributes, such as birth years and population numbers, are encoded as monotonic representations in the activation space of LLMs and can be manipulated with interventions. In turn, Stolfo et al. (2023) showed that, in autoregressive LMs, the operations and numerical information necessary for solving quantitative reasoning are processed in the lower layers of the model, and these results are used by the attention layers to predict the final calculation outcomes. Zhu et al. (2025) studied the representation of numbers in language models’ hidden states during single-hop arithmetic tasks (e.g., What is the sum of 12 and 34?). Their analysis revealed that numerical information is encoded linearly within the hidden states and demonstrated that individual digits could be manipulated independently. In this study, we add to this literature by introducing incremental arithmetic problem solving, i.e., what numerical information is contained in the model’s hidden states at each time step of multi-hop arithmetic reasoning.
Model interpretability methods.
Linear probing Alain and Bengio (2017b) is one of the representative methods for analyzing the internal representations of neural models—a small model predicts a specific feature from them, thereby determining whether the input contains information about that feature. In this study, we use them to derive the models’ intermediate answers. The causality with the model’s output can be further verified by examining if a model’s predictions change when the hidden states are intervened Li et al. (2023); Wu et al. (2023b). One representative intervention method is activation patching Vig et al. (2020); Meng et al. (2022); Zhang and Nanda (2024), where hidden states obtained from one model instance are transplanted onto another during inference to change its predictions. Such techniques can be applied as a way to control model behavior in practical scenarios such as mitigating inherent biases Zhao et al. (2019); Ganguli et al. (2023); Yang et al. (2024b). Here, we employ activation patching to validate the plausibility of the probing results.
@@ 3.1 Arithmetic problems
We prepared a dataset of multi-hop arithmetic problems similarly to Kudo et al. (2023) and Yu (2025). Each sample is a string of assignments (e.g., A=1) and operations (e.g., B=1+3 or B=1+A) ending with a query for a variable’s value (e.g., B=?). We also defined five complexity levels, depending on (i) how many equations need to be resolved to reach the answer (#Step in Table 1), (ii) how many variables’ values cannot be immediately resolved (and thus pended to a stack) in their first appearance when incrementally reading the problem from left to right (#Stack), (iii) and the number of unnecessary distractor equations (#Dist.). For example, in the Level 5 example in Table 1, where #Step is three and #Stack is two, calculating A requires at least three steps of reasoning: C(=1+2)=3, B(=2+3)=5, and then A(=1+5)=6, and two variables need to be resolved before reaching A: B and C.
@@
Formally, let $v$ denote a variable name sampled from the 26 letters of the English alphabet $\Sigma=\{\text{a},\text{b},\text{c},\dots,\text{z}\}$, and $d$ a number sampled from the set of decimal digits $D=\{0,1,2,\ldots,9\}$. Each instance consists of multiple equations $e_{1},e_{2},\cdots,e_{n}]$ followed by a final query $q$. Each equation follows the format $v=d$, $v=d\pm d$, or $v=d\pm v$. We denote $i$-th variable to appear within an instance from the left as $v_{i}$. E.g., in the Level 5 example in Table 1, $v_{1}=\texttt{A}$, $v_{2}=\texttt{B}$, and $v_{3}=\texttt{C}$. [^2] For ease of reading, all examples throughout this paper use the uppercased variables A, B, C, and only the operator +. The value assigned to a variable $v_{i}$ is denoted as $\{v_{i}\}\in D$.
Generation rules.
We ensure that $\{v_{i}\}$ for any $v_{i}$ is also a single-digit number, and $\{v_{i}\}$ is constant within the same instance (i.e., we exclude cases such as A=1+2,A=B+2,B=6). All samples of the same complexity level follow the exact same format except for the actual numbers, variable names, and operators, meaning that $\{v_{i}\}$ in each level can be obtained with exactly the same abstract procedure. E.g., first calculate $\{v_{3}\}$, and then calculate $\{v_{2}\}$ with $\{v_{3}\}$, and then finally calculate $\{v_{1}\}$ with $\{v_{2}\}$. Non-duplicated instances are created for each level by varying the variable names, numbers, and operators appearing in the equations. We created 12,000 unique instances in total, of which 10,000 are used for training probing classifiers and 2,000 for testing their accuracy. To prevent the probe itself from performing arithmetic, we constructed the training and test sets so that no duplicate arithmetic expressions appear. [^3] For example, if 1+2 appears in the training set, then 1+2 does not appear in the test set.
@@ 3.2 CoT-style inference
Henceforth, we refer to the part before CoT (explained in 3.1) as the Input and the CoT-reasoning part as the Output of an instance. The Output, as shown in the right part of Table 1, is also a sequence of equations in the same format as the Input and can be split into intermediate steps $z$ and a final answer $y$. For example, for the Level 1 instance in Table 1 (topmost), x=x=A=1+B, B=2,” z=z=A=1+B, B=2, A=1+2, A=,” and y=y=3.” That is, the task is to generate intermediate steps $z$ and derive a final answer $y$ given an Input $x$ and a demonstration of three examples. As a sanity check, we confirmed that the target models could follow the expected Output format and solve the tasks with nearly 100% accuracy in this setting (see §˜A.1 for more details).
@@
We denote a token position within the entire concatenated sequence $x\oplus z\oplus y$ with $\in\mathbb{Z}. tt$ is relative to CoT; that is, $t$ is zero where CoT begins, negative within the Input, and positive within the Output. Similarly, we assign an equation position $t_{\mathrm{eq}}\in\mathbb{Z}$ to each equation in the Input and Output (subscripts on the underlines in Table 1).
4 Probing
When do models solve (sub)problems in CoT-style reasoning? Do they (i) finalize reasoning during the Input stage, with the CoT as a post-hoc explanation (Think-to-Talk), or (ii) solve the task step-by-step during CoT generation (Talk-to-Think)? We address this by examining where the final answer, or the necessary information for it, emerges in the model’s internal representations using linear probes.
4.1 Training settings for linear probes
@@
We train a linear probe Alain and Bengio (2017a) for an $l$-layer LLM ($l=0$ for the input embedding layer). To identify where the LLM solved a particular (sub)problem, we train a separate probe for each combination of token position $t\in\mathbb{Z}$, layer depth $l\in\mathbb{N}$, and $v_{i}\in\Sigma$ in each level of the problem. Specifically, given a model’s $d$-dimentional hidden state $\bm{h}_{t,l}\in\mathbb{R}^{d}$, the probing classifier $f_{t,l,v_{i}}(\cdot):\mathbb{R}^{d}\to D$ predicts $\{v_{i}\}$. That is, for each $(t,l)$, we first obtained 10,000 of $\bm{h}_{t,l}$ from training instances and then evaluated the accuracy of $f_{t,l,v_{i}}(\cdot)$ for each $v_{i}$ with 2,000 hidden states from test instances and the correct $\{v_{i}\}$. If a probe $f_{t,l,v_{i}}(\cdot)$ achieves high accuracy, this suggests that $\{v_{i}\}$ is already computed at the corresponding position ($t$ and $l$). Figure 2 illustrates the probing results with a Level 3 task, where, for example, the value of B can be extracted within Input and thus already computed before CoT begins.
The probe $f_{t,l,v_{i}}(\cdot)$ is a single linear transformation; that is, the probe is applied to $\bm{h}_{t,l}$ as follows:
$\hat{\$\{v_{i}\}}_{t,l,}$ $=f_{t,l,v_{i}}(\bm{h}_{t,l})$ (1)
@@
where, $\bm{W}_{t,l,v_{i}}\in\mathbb{R}^{|D|\times d}$ and $\bm{b}_{t,l,v_{i}}\in\mathbb{R}^{|D|}$ are the weight and bias parameters of the probe, respectively. The symbol $\hat{\cdot}$ is used to refer to the model’s estimate. We train the probes using stochastic gradient descent Robbins (1951) to minimize the cross entropy loss. The hyperparameters are listed in Table 5 in the appendix.
Figure 2: Probing results for Qwen2.5-7B at the task Level 3. The heatmaps in the lower section represent the accuracy of probes computed on the evaluation set. Each cell shows the probing accuracies in each token $t$, layer $l$. The upper part indicates the maximum probing accuracy achieved at each token position $t$. The input sequence below the line graphs is just an example; in the actual evaluation set, each variable name, number, and operator are randomly sampled from $(D,\Sigma,\{+,-\})$.

### v4

# LLMs Faithfully and Iteratively Compute Answers During CoT: A Systematic Analysis With Multi-step Arithmetics

@@ Abstract
This study investigates the internal information flow of large language models (LLMs) while performing chain-of-thought (CoT) style reasoning. Specifically, with a particular interest in the faithfulness of the CoT explanation to LLMs’ final answer, we explore (i) when the LLMs’ answer is (pre)determined, especially before the CoT begins or after, and (ii) how strongly the information from CoT specifically has a causal effect on the final answer. Our experiments with controlled arithmetic tasks reveal a systematic internal reasoning mechanism of LLMs. They have not derived an answer at the moment when input was fed into the model. Instead, they compute (sub-)answers while generating the reasoning chain on the fly. Therefore, the generated reasoning chains can be regarded as faithful reflections of the model’s internal computation.
LLMs Faithfully and Iteratively Compute Answers During CoT:
A Systematic Analysis With Multi-step Arithmetics
Keito Kudo1,2, Yoichi Aoki1,2, Tatsuki Kuribayashi3,1, Shusaku Sone1 Masaya Taniguchi2,1, Ana Brassard2, Keisuke Sakaguchi1,2, Kentaro Inui3,1,2 1Tohoku University, 2RIKEN, 3MBZUAI {keito.kudo.q4, youichi.aoki.p2, sone.shusaku.r8}@dc.tohoku.ac.jp, {tatsuki.kuribayashi, kentaro.inui}@mbzuai.ac.ae, keisuke.sakaguchi@tohoku.ac.jp, {masaya.taniguchi, ana.brassard}@riken.jp
@@ 1 Introduction
Modern large language models (LLMs), given a query from users, typically produce intermediate reasoning steps as well as a final answer. Such reasoning steps to achieve the answer, i.e., explanation, are helpful to grasp the model’s underlying reasoning process and the plausible justification for the produced answer. One critical concern is how faithful the explanation, typically with a chain-of-thought (CoT) Wei et al. (2022) style, is to their final answer, i.e., the causal relationship between the CoT process and the final answer. In particular, how models internally associate the CoT part and its following answer is an intriguing question in the interpretability field. In this study, we analyze the faithfulness between the CoT explanation and the model’s final answer through the lens of (mechanistic) interpretability. This question involves two key subquestions—when, in the first place, models come up with (sub-)answers during CoT-style reasoning, and how these are referred to internally. One possible strategy, for example, would be to reach the final answer even during reading problem statements before CoT generation (the first pass), and then the model produces a fluent explanation just to follow an expected format, and ultimately, the model simply refers to their predetermined answers. In this case, the final answer is no longer strictly faithful to the CoT explanation, and users should be aware of this limitation in interpreting the model’s outputs.
@@
Figure 1: Using linear probes, we investigated at which time during the LLM’s problem-solving process it is possible to determine the values of each variable, illustrating the model’s problem-solving process. Our analysis indicates that LLMs come up with (sub-)answers during CoT. This conclusion is also consistent with the findings from the causal experiments in §˜4.
@@
In our first experiments, we exploratorily analyze when (at each layer at each timestep) models represent answers internally, using linear probes. To prevent situations where CoT is not required for deriving the answer, we set up controlled testbeds of symbolic arithmetic reasoning and observed when/where probes could accurately extract and control the answer (Figure 1). By comparing accuracies across each timestep, one can observe at which point models internally start being informative (the values of the variables are encoded in the hidden states as linearly separable representations) to the probes, illustrating the model’s internal reasoning pattern. We found that models have not derived an answer at the moment they first read the problem; instead, they obtain (sub-)answers while generating the reasoning chain, suggesting the faithfulness of CoT to the final answer.
Based on the probing results, we further conducted causal intervention analyses to clarify the causal relationship between the model’s internal representations and answers (§˜4). We found that, after CoT generation, the model’s final answer can be changed by causal intervention to internal representations for the CoT explanation part, while not for the first pass part. The discovered causal graph follows recency bias, i.e., each reasoning step in CoT causally depends on its previous step, and the final answer heavily relies on the latter part of CoT. Thus, we tentatively conclude that models derive an answer during CoT on the fly, and the generated reasoning chains can be regarded as faithful reflections of the model’s final answer.  [^1] Code and data are available at https://github.com/keitokudo/faithful-cot-multistep-arithmetic
@@
Table 1: Examples of arithmetic reasoning tasks used in our experiments at each complexity level. #Step indicates the number of required operations to reach the final answer. #Stack indicates how many variables’ values are not immediately determined in their first appearing equation. #Dist. is the number of unnecessary distractor equations. The number (e.g., $\mathbf{-3}$) indicated in the lower right corner of each equation represents the equation’s position. This position is used as a reference point for calculating $t^{*}_{\mathrm{eq}}$ in §˜3.2.
@@ 2.1 Arithmetic problems
We prepared a synthetic dataset of multi-hop arithmetic problems similarly to Kudo et al. (2023) and Yu (2025). Each problem instance consists of strings of assignments (e.g., A=1) and operations (e.g., B=1+3 or B=1+A) ending with a query for a variable’s value (e.g., B=?). By using this synthetic dataset, we can conduct controlled experiments across different levels of problem complexity and confirm the generality of the obtained results. Moreover, by training and evaluating probes for each common problem format (level), we can conduct fine-grained, token-level analyses. Such controlled analyses are difficult with datasets written in natural language.
We specifically defined five complexity levels (Table 1), depending on (i) how many equations need to be resolved to reach the answer (#Step in Table 1), (ii) how many variables’ values cannot be immediately resolved (and thus pended to a stack) in their first appearance when incrementally reading the problem from left to right (#Stack), (iii) and the number of unnecessary distractor equations (#Dist.). For example, in the level 5 example in Table 1, where #Step is three and #Stack is two, calculating A requires at least three steps of reasoning: C(=1+2)=3, B(=2+3)=5, and then A(=1+5)=6, and two variables need to be resolved before reaching A: B and C.
@@
Formally, let $v$ denote a variable name sampled from the 26 letters of the English alphabet $\Sigma=\{\text{a},\text{b},\text{c},\dots,\text{z}\}$, and $d$ denote a number sampled from the set of decimal digits $D=\{0,1,2,\ldots,9\}$. Each instance consists of multiple equations $e_{1},e_{2},\cdots,e_{n}]$ followed by a final query $q$. Each equation follows the format $v=d$, $v=d\pm d$, or $v=d\pm v$. We denote $i$-th variable in an instance from the left as $v_{i}$; e.g., in the Level 5 example in Table 1, $v_{1}=\texttt{A}$, $v_{2}=\texttt{B}$, and $v_{3}=\texttt{C}$. [^2] For brevity, all examples in this paper use the uppercased variables A, B, C, and only the operator +, but the actual instances have a variety in the variable names and operator types. The value assigned to a variable $v_{i}$ is denoted as $\mathdollar\{v_{i}\}\in D$.
Data constraints.
We ensure that $\mathdollar\{v_{i}\}$ for any $v_{i}$ is also a single-digit number, and $\mathdollar\{v_{i}\}$ is constant within the same instance (i.e., we exclude cases such as A=1+2,A=B+2,B=6). All samples of the same complexity level follow the exact same format except for the actual numbers, variable names, and operators. In other words, an answer $\mathdollar\{v_{i}\}$ in each level can be obtained with exactly the same procedure. For example, first calculate $\mathdollar\{v_{3}\}$, and then calculate $\mathdollar\{v_{2}\}$ with $\mathdollar\{v_{3}\}$, and then finally calculate $\mathdollar\{v_{1}\}$ with $\mathdollar\{v_{2}\}$. Non-duplicated instances are created for each level by varying the variable names, numbers, and operators appearing in the equations. We use 10,000 instances to train probing classifiers and 2,000 to test their accuracy. To prevent the probe from simply memorizing (sub-)answers, the training and test sets were constructed with no overlap in arithmetic expressions. [^3] The overlap is computed at the equation-level. For example, if 1+2 appears in the training set, then 1+2 does not appear in any test instances.
@@ 2.2 CoT-style inference
Given the input described in §˜2.1, the model generates CoT $z$ and a final answer $y$. Henceforth, we refer to the part before CoT as the Input $x$ and the CoT-reasoning part ($y,z$) as the Output of an instance, as shown in the right part of Table 1. For example, for the Level 1 instance in Table 1 (topmost), x=x=A=1+B, B=2,” z=z=A=1+B, B=2, A=1+2, A=”, and y=y=3.” We demonstrate three examples in the same problem level for the model to inform the problem setting and expected CoT-style reasoning style (Table 1). That is, the task is, given a demonstration of three examples, and an Input $x$, to generate intermediate steps $z$ and derive a final answer $y$. As a sanity check, we confirmed that the target models could follow the expected Output format and solve the tasks with nearly 100% accuracy in this setting (§˜B.1).
2.3 Research scope
We are particularly concerned about the faithfulness between the CoT part $z$ and its following final answer $y$, since these two parts are typically shown to users, and thus are of primary concern regarding the model’s reliability from the user’s perspective. In contrast, some recent studies Afzal et al. (2025); Cox (2025); Ye et al. (2025) have investigated whether the CoT $z$ depends on the internal state of the model $x$ before CoT begins (whether the CoT $z$ is post hoc or not). This causal relationship $x\rightarrow z$ is out of scope in this study, while we are still interested in whether the model $x$ has a predetermined answer $y$ in order to distinguish the possibilities of $x\rightarrow y$ and $z\rightarrow y$, i.e., which part is causally relied on to the final answer $y$.
3 Linear probing
To begin with the analysis, it will be good to know where the final answer, or the necessary sub-answer for it, can possibly come from. We first analyze where final-/sub- answers can be extracted from the model’s internal representations, using linear probes.
Position $t$.
@@ 3 Linear probing
We denote a token position within the entire concatenated sequence $x\oplus z\oplus y$ with $t\in\mathbb{Z}$. The position $t$ is relative to CoT; that is, $t$ is zero at the moment CoT begins, negative within the Input, and positive within the Output. Similarly, we assign an equation position $t_{\mathrm{eq}}\in\mathbb{Z}$ to each equation in the Input and Output (subscripts on the underlines in Table 1).
3.1 Training linear probes
@@
Linear probes Alain and Bengio (2017a) are trained to identify when the model derives the answer internally. We separately train a probe for each combination of token position $t\in\mathbb{Z}$, layer depth $l\in\mathbb{N}$, and the variable of interest $v_{i}\in\Sigma$ in each level of the problem. That is, each probe is solely responsible for the position $(t,l,v_{i})$ to identify the answer $\hat{\mathdollar\{v_{i}\}}$ represented. Formally, given a model’s $d$-dimensional hidden state $\bm{h}_{t,l}\in\mathbb{R}^{d}$, the probing classifier $f_{t,l,v_{i}}(\cdot):\mathbb{R}^{d}\to D$ is trained to predict the correct answer $\mathdollar\{v_{i}\}$. If a probe $f_{t,l,v_{i}}$ achieves high accuracy, this suggests that the answer for a particular variable $\mathdollar\{v_{i}\}$ is already computed at the corresponding position $(t,l)$. The probe $f_{t,l,v_{i}}$ is a single linear transformation:
$\hat{\mathdollar\{v_{i}\}}_{t,l}$ $=f_{t,l,v_{i}}(\bm{h}_{t,l})$ (1)
@@
where, $\bm{W}_{t,l,v_{i}}\in\mathbb{R}^{|D|\times d}$ and $\bm{b}_{t,l,v_{i}}\in\mathbb{R}^{|D|}$ are the trainable weight and bias parameters of the probe, respectively. The symbol $\hat{\cdot}$ refers to the model’s predicted answer. We train the probes using stochastic gradient descent Robbins (1951) to minimize the cross entropy loss. The hyperparameters are listed in Table 8 in the appendix.
Each probe classifier is trained with 10,000 of $\bm{h}_{t,l}$ from training instances and then computes the accuracy with 2,000 hidden states from test instances.
Figure 2: Probing results for Qwen2.5-7B at the Level 3 task. The heatmaps in the lower section represent the accuracy of probes computed on the evaluation set. Each cell shows the probing accuracies in each token $t$, layer $l$. The upper part indicates the maximum probing accuracy achieved at each token position $t$. The input sequence below the line graphs is just an example; in the actual evaluation set, each variable name, number, and operator are randomly sampled from $(D,\Sigma,\{+,-\})$.